Problem 15.16 Assuming the rings are planar, label each ion as aromatic, antiaromatic, or not aromatic a. D. b. D:-
Answer #1

44+2 = 2 4h - 2-2=0 n=0 Aromatic 4 h = 4 n=1 Antiaromatic 4n=0 n=2 Antiaromatic 4h+2=10 un = 10-2=8 hn=8 n=2. frornatic